fentyf4iry:

Volume of cylinders 101 with Fentyf4iry (●'◡'●) *Please don't comment until I am done!*

fentyf4iry:

Let's start out with a word problem, first read the question and write the radius and height down in a piece of paper so you understand what to do next.

1 attachment
fentyf4iry:

Radius = 17 Height = 19

fentyf4iry:

Use these numbers in the volume formula. Use 3.14 for ​𝜋 (pi).

fentyf4iry:

Volume = ​𝜋r²h ≈ 3.14 × 17× 17 ×19 ≈ 17,241.74

fentyf4iry:

The volume of the cylinder is about 17,241.74 cubic meters.

fentyf4iry:

This will be your total if you do it in a calculator!

1 attachment
Extrinix:

Just for volumes? might as well do surface area as well \(2\pi r^2+2\pi rh\)

fentyf4iry:

@extrinix wrote:
Just for volumes? might as well do surface area as well \(2\pi r^2+2\pi rh\)
i'll do a surface area tutorial soon!

fentyf4iry:

*CLOSED*